OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 2x + 4 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

Find D: \(D = b^2 - 4ac\) D = 4 - 16 = -12 So, you will get no real solutions..

OpenStudy (anonymous):

\[\huge D < 0 \implies No \quad \mathbb{R} \quad Solutions\]

OpenStudy (angela210793):